# Learning General Representation of 12-Lead Electrocardiogram With a Joint-Embedding Predictive Architecture

### v1

@@ 6.2 Fine-tuning
Fine-tuning is another method to evaluate the quality of learned representations, as it tests the model’s ability to adapt its pre-trained features to new tasks. In the fine-tuning process, we add a linear classification head at the end of the encoder and train the whole network for 10 epochs. Fine-tuning can potentially enhance performance beyond what is achieved with linear evaluation alone.
Table 3 presents the results of fine-tuning on the PTB-XL and CPSC2018 datasets. ECG-JEPA is compared with other SSL methods in a multi-class classification setting. The results indicate that ECG-JEPA achieves the highest AUC and F1 scores on PTB-XL and the highest AUC on CPSC2018. These results highlight ECG-JEPA’s effectiveness in fine-tuning scenarios, demonstrating its ability to leverage transfer learning by refining and adapting its learned representations for improved performance in specific classification tasks.
@@
MoCo v3111Score as reported in [23]. 800 0.913 0.644 0.967 0.838
MTAE [^1] Score as reported in [23]. 800 0.910 0.613 0.961 0.769
MLAE [^1] Score as reported in [23]. 800 0.915 0.625 0.973 0.816
CMSC [^1] Score as reported in [23]. 800 0.877 0.510 0.938 0.717
@@ 6.4 Reduced Lead Evaluation
Because transformer architectures can accept variable input lengths, we can evaluate ECG-JEPA’s performance on reduced lead tasks. In this experiment, we perform a linear evaluation on the PTB-XL multi-label task using reduced leads. Specifically, we utilize only a single lead (Lead II) and two leads (Lead II and V1), training linear classifiers on top of the learned representations for 10 epochs. We then compare ECG-JEPA’s performance with ST-MEM, another model supporting reduced lead evaluation. Table 5 presents the results of this reduced lead linear evaluation. Notably, ECG-JEPA performs well even in the reduced lead setting. This robustness is particularly valuable since most daily mobile devices typically output only one or two leads, highlighting ECG-JEPA’s potential for practical applications in mobile health monitoring.
Table 5: Reduced lead evaluation. Linear evaluation on PTB-XL multi-label task with reduced lead settings.
@@ 6.5 ECG Feature Extraction
ECG features are essential for diagnosing and monitoring cardiac conditions. We evaluate the model’s ability to extract key ECG features, including heart rate and QRS duration, from the learned representations of the PTB-XL dataset. Unlike classification tasks, which are highly perceptual, ECG features are directly related to the morphology of the signal. For example, heart rate can be measured by the average distance between R-peaks, and QRS duration can be measured by the average width of the QRS complex.
Several methods have been proposed to segment ECG signals [30, 31, 32, 33]. Segmentation models can be used to extract ECG features. In this experiment, we employed a publicly available segmentation model [33] to extract ECG features from the PTB-XL dataset, which served as our ground truth labels. We then trained a linear regression model on the learned representations to predict these ECG features, using the mean squared error (MSE) as the loss function.
Table 6 presents the performance comparison, showing the means and standard deviations of the absolute differences between the extracted and predicted values across all samples in the PTB-XL test set.
Interestingly, although the model’s learned representations are implied to capture high-level features based on their strong performance in linear evaluation tasks across various datasets, they still retain the ability to recover low-level information, such as ECG features. This dual capability of capturing both high-level and low-level information is noteworthy and highlights the model’s versatility.
@@ 7.1 Effect of CroPA
Table 7 presents the results of our evaluation of the effectiveness of CroPA. CroPA introduces a “human-like" inductive bias, enabling the model to be trained more efficiently on multi-lead ECG data. Without CroPA, models may require more epochs to converge. For a fair comparison, we trained ECG-JEPA with and without CroPA for 100 and 200 epochs and compared their performance on the PTB-XL multi-class task. The results show that CroPA considerably improves the model’s performance, demonstrating its effectiveness in capturing inter-lead relationships and enhancing the model’s ability to learn meaningful representations.
@@ 9 Conclusion
Appendix A Supplementary Materials
A.1 Experimental Details on ECG-JEPA
@@
A.2 Pretraining Other SSL Models
@@ A.3 Downstream Datasets Details
Table 12, 13, 14, and 15 show the distribution of the PTB-XL and CPSC2018 datasets in both multi-label and multi-class settings.
@@
Table 12: PTB-XL multi-label distribution. The sum of samples in each class exceeds the total number of ECG recordings due to the multi-label nature of the dataset.
@@
Table 13: PTB-XL multi-class distribution. ECGs with multiple labels are excluded from the PTB-XL multi-label dataset.
@@
Table 14: CPSC2018 multi-label distribution. The sum of samples in each class exceeds the total number of ECG recordings due to the multi-label nature of the dataset.
@@
Table 15: CPSC2018 multi-class distribution. The number of ECG recordings for each superclass in the PTB-XL dataset. Note that the sum of samples in each class exceeds the total number of ECG recordings due to multi-label the nature of the dataset.
@@
A.4 Exponential Moving Average
@@
A.5 Sufficiency of 8 Leads in ECG Analysis
@@
A.6 Comparison with 12-Lead Model

### v2

@@ 6.2 Fine-tuning
Fine-tuning is another method to evaluate the quality of learned representations, as it tests the model’s ability to adapt its pre-trained features to new tasks. In the fine-tuning process, we add a linear classification head at the end of the encoder and train the entire network for 10 epochs. Similar to linear evaluation, training for 10 epochs is sufficient, as further training does not lead to additional performance gains. Fine-tuning can potentially enhance performance beyond what is achieved with linear evaluation alone.
Table 3 presents the results of fine-tuning on the PTB-XL and CPSC2018 datasets. ECG-JEPA is compared with other SSL methods as well as supervised methods in a multi-class classification setting, where the student network is trained directly for the supervised task. The results indicate that ECG-JEPA achieves the highest AUC and F1 scores on PTB-XL and the highest AUC on CPSC2018. These findings underscore ECG-JEPA’s effectiveness in fine-tuning scenarios, demonstrating its ability to refine and adapt learned representations for improved performance in specific classification tasks.
@@
Supervised 100 0.887 0.608 0.893 0.566
MoCo v3111 800 0.913 0.644 0.967 0.838
MTAE111 800 0.910 0.613 0.961 0.769
MLAE111 800 0.915 0.625 0.973 0.816
CMSC111 800 0.877 0.510 0.938 0.717
@@ 6.4 Reduced Lead Evaluation
Since transformer architectures can handle variable input lengths, we evaluated ECG-JEPA’s performance with reduced leads. In this experiment, we conducted a linear evaluation on the PTB-XL multi-label task using only a single lead (Lead II) and two leads (Lead II and V1), training linear classifiers on the learned representations for 10 epochs [^3] We compare only with ST-MEM, as it is a transformer-based model whose pretrained weights are publicly available.. Table 5 presents the results. Notably, ECG-JEPA maintains strong performance even with fewer leads, which is valuable for practical applications in mobile health monitoring, where most devices typically output only one or two leads.
Table 5: Reduced lead evaluation. Linear evaluation of PTB-XL multi-label classification in single-leade (II) and dual-lead (II and V1) configurations.
@@ 6.5 ECG Feature Extraction
Extracting ECG features is crucial for diagnosing and monitoring cardiac conditions. In this experiment, we assess the model’s ability to extract key features such as heart rate and QRS duration from the learned representations of the PTB-XL dataset. Unlike classification tasks, which focus on perceptual patterns, ECG features are directly tied to the signal’s morphology. For example, heart rate is derived from the average distance between R-peaks, while QRS duration is measured by the average width of the QRS complex.
Various methods exist for segmenting ECG signals [28, 29, 30, 31], which can be used to extract ECG features. For this experiment, we utilized a publicly available segmentation model [31] to generate ground truth labels for heart rate and QRS duration from the PTB-XL dataset. We then trained a linear regression model on the learned representations to predict these features, using mean squared error (MSE) as the loss function.
Table 6 shows the performance comparison, reporting the means and standard deviations of the absolute differences between the predicted and extracted values for the heart rate and QRS duration across the PTB-XL test set.
Interestingly, although the model’s representations are designed to capture high-level features, they retain the capacity to recover low-level ECG features. This dual ability to encode both high-level semantics and low-level morphology underscores the versatility of ECG-JEPA, highlighting its potential in both diagnostic and real-world applications.
@@ 7.1 Effect of CroPA
Table 7 presents the results of our evaluation of the effectiveness of CroPA. CroPA introduces a “human-like" inductive bias, enabling the model to be trained more efficiently on multi-lead ECG data. Without CroPA, models may require more epochs to converge. For a fair comparison, we trained ECG-JEPA with and without CroPA for 100 and 200 epochs and compared their performance on the PTB-XL multi-class task. The results show that CroPA improves the model’s performance, demonstrating its effectiveness in capturing inter-lead relationships and enhancing the model’s ability to learn meaningful representations.
@@ 9 Conclusion
10 Data Availability
All datasets used in this work are publicly available, and our code can be accessed at https://github.com/sehunfromdaegu/ecg_jepa.
11 Funding
This work was supported by the National Research Foundation of Korea(NRF) Grant 2023R1A2C1005562, funded by the Korean government.
Appendix A Experimental Details
A.1 Hyperparameters for ECG-JEPA
@@
A.2 Hyperparameters for Other Pretrained Models
@@ A.3 Downstream Datasets Details
Table 12, 13, 14, and 15 show the distribution of the PTB-XL and CPSC2018 datasets in both multi-label and multi-class settings. Note that the sum of samples in each class exceeds the total number of ECG recordings in multi-label task.
@@
Table 12: PTB-XL multi-label distribution.
@@
Table 13: PTB-XL multi-class distribution.
@@
Table 14: CPSC2018 multi-label distribution.
@@
Table 15: CPSC2018 multi-class distribution.
@@
Appendix B Exponential Moving Average
@@
Appendix C Sufficiency of 8 Leads in ECG Analysis
@@
Appendix D Comparison with 12-Lead Model